OpenStudy (gabylovesyou):

Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. If Michelle returned the movie and paid $4.00, how many nights did she keep the movie?

hartnn (hartnn):

for 1st night she paid $1.50 for 2nd night total she paid = 1.50+1.25 = $2.75 for 3rd night total she paid =2.75+1.25 =$4 so she kept it for 3 days

OpenStudy (gabylovesyou):

thank you :)

hartnn (hartnn):

welcome :)